OpenStudy (anonymous):

i will give a medal Find the value for x that makes the equation 1.5x+6=3(2.5x-6) true. plzz show me how to do this

OpenStudy (anonymous):

Firstly, we need to expand the brackets on the right. Can you do that for me?

OpenStudy (anonymous):

1.5x+6=3x-15 @tom982

OpenStudy (anonymous):

Not quite. \[3(2.5x-6)=3*2.5x-3*6\] Do you see why it's wrong?

OpenStudy (anonymous):

ohhhh distributive property?

OpenStudy (anonymous):

@tom982

OpenStudy (anonymous):

Yep :)

OpenStudy (anonymous):

1.5x + 6 = 7.5x - 18 (we multiply the values inside the brackets) 6 + 18 = 7.5x - 1.5x (we transfer the 18 on the other side and 1.5x on the one) 24 = 6x (sum up both numbers) 4 = x (divide) This should help :)

OpenStudy (anonymous):

ooops wrong thing

OpenStudy (anonymous):

i wrote it wrong

OpenStudy (anonymous):

First u need to expand the bracket. second shif all the x to the left hand side. Then divide both sides. \[1.5x+6=3(2.5x-6)\]

OpenStudy (anonymous):

\[1.5x+6=7.5x-18\]

OpenStudy (anonymous):

\[1.5x-7.5x=-18-6\]

OpenStudy (anonymous):

\[-6x=-24\]

OpenStudy (anonymous):

Then divide -6 on both sides.

OpenStudy (anonymous):

\[\frac{ -6 }{ -6 }x=\frac{ -24 }{ -6 }\]

OpenStudy (anonymous):

\[x=?\]solve it @tarap0123

OpenStudy (anonymous):

4 right?

OpenStudy (anonymous):

@MARC_

OpenStudy (anonymous):

great

OpenStudy (anonymous):

thank you soooo much @MARC_

OpenStudy (anonymous):

and also thank you @tom982

OpenStudy (anonymous):

yw! @tarap0123

OpenStudy (anonymous):

My pleasure.